OpenStudy (anonymous):

Is the following equation always true, sometimes true, or never true? 1) x + x = 2x 2) x times x = 2x 3) x^2 = 9 (sometimes?)

OpenStudy (anonymous):

x+x=2x

OpenStudy (anonymous):

x times x = x sqaured

OpenStudy (anonymous):

*squared

OpenStudy (anonymous):

Never true!

OpenStudy (anonymous):

1 and 2 are always true but 3 is sometimes

OpenStudy (anonymous):

IT DEPENDS UPON THE CONDITION WHEREVER IT IS GIVEN OR NEEDED. BTW THE FIRST ONE IS ALWAYS CORRECT AS I REMMEMBER.

OpenStudy (paxpolaris):

1) \(x + x = 2x\) always 2) \(x \times x = 2x\) sometimes ... when \(x=2\) 3) \(x^2 = 9\) sometimes ... when \(x=3\), or \(x=-3\)

OpenStudy (anonymous):

@PaxPolaris IS RIGHT.